0 0
Add a comment Improve this question Transcribed image text
Answer #1

a)
Magnetic lines are coming into the paper
=> View faces Magnetic North
Since magnetic lines are from North to south

But Magnetic North is Goegraphical South
Hence Answer: B

b)
Flux=B.A=BA cos(theta)
where A is area vector whose magnitude is pie*R^2 and direction is perpendicular to ring
In this view theta=0
=> Flux=BA
5.8*10^-3=5*10^-5*pie*R^2
=> R=6.0765m

c)
As the loop is rotating Flux through it will start decreasing since theta increases

=> According to lenz's law i will be in that direction in which change is opposed
=> Current will be clockwise

d) Flux is zero when angle between B and A is 90
=> loop has rotaed to 90 degrees
option (B)

e)
emf developed in ciruit: change in flux/time taken for that change
=5.8*10^-3/4=1.45mV
Power=V^2/R =(1.45*10^-3)^2/(3.5*10^-5)
=0.06W

f)
answer: B